OpenStudy (anonymous):

what does it mean by reflect the triangle over x=-2

OpenStudy (amistre64):

what does a reflection do?

OpenStudy (anonymous):

It flips over a line

OpenStudy (amistre64):

correct, so flip it over the vertical line that is marked at x=-2

OpenStudy (amistre64):



OpenStudy (amistre64):

spose we had a point (-6,3) move it to add 2 to the x parts to move it and flip about the x=-2+2 axis (-4,3), now flip (4,3), now move it back again by shifting it to the x=0-2 (2,3) or do you know how to determine a midpoint?

OpenStudy (amistre64):

(x1 + x2)/2 = -2 (-6 + x2)/2 = -2 -6 + x2 = -4 x2 = 2

OpenStudy (anonymous):

Still don't get it

OpenStudy (amistre64):

 the distance from the line x=-2 remains the same after you flip it ....

OpenStudy (amistre64):

therefore -2 is the middle point between your required x points